OpenStudy (anonymous):

a bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking a black marble?

OpenStudy (anonymous):

P(R and B) = P(R) * P(B)

OpenStudy (anonymous):

Can u find P(R) ?

OpenStudy (anonymous):

uhh...i dont exactly know what P(R) is.

OpenStudy (anonymous):

Probability of picking a red marble

OpenStudy (anonymous):

5/15 so 1/3

OpenStudy (anonymous):

Yes. Now find P(B)

OpenStudy (anonymous):

4/15

OpenStudy (anonymous):

Great Now, find P(R and B)

OpenStudy (anonymous):

9/15...3/5

OpenStudy (anonymous):

P(R and B) =P(R) * P(B) = 1/3 *4/15 =?

OpenStudy (anonymous):

4/45 ? i think

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

welcome